# Enabling Robust Exoplanet Atmospheric Retrievals with Gaussian Processes

### v1

@@ 3.2 Recovery of Injected Noise and Feature
In all three cases, we find that the retrieval frameworks are able to provide both precise and accurate estimates of the underlying correlation. The median retrieved values of all hyperparameters in each case are within of the true values, and all hyperparameters are constrained. Although the local kernel does appear to slightly overestimate the wavelength and width of the injected feature in both the local-only and combined case, these do not significantly affect its ability to retrieve accurate posterior distributions for atmospheric parameters, and are still within a reasonable margin of error away from the injected values. Specifically, the retrieved central wavelength is only 0.01 m from the true value, which we attribute to randomness in the noise instance used to simulate the data and the low resolution of the data. The exact central wavelength may be difficult to constrain for a feature characterized by only 5 data points, but 0.01 m is within a single bin width of the true wavelength. As such, we consider this an accurate inference.
Notably, the global-only case retrieves a higher global correlation amplitude than the combined case. This is due to the absence of a local kernel in the global-only case, leading to overcompensation; to account for the 1.65 m feature without a local kernel, the global kernel retrieves a higher amplitude that can marginalize over the feature. Thus, in the combined case, the inferred amplitude is lower, with the true injected value being within of the median inferred amplitude. The GP contribution functions for all four cases are shown in Appendix B.
@@ 4 Application to the JWST NIRISS/SOSS Spectrum of WASP-96 b
Radica et al. (2023) presented the near-infrared (NIR) transmission spectrum of WASP-96 b observed with JWST NIRISS/SOSS (0.6-2.8 m) as part of the JWST Early Release Observations (ERO) program. The spectrum is split into two overlapping SOSS orders (0.6–1.0 m and 0.85–2.8 m), with an additional third order not considered within the data reduction process. The data was reduced via the exoTEDRF pipeline111Formerly known as supreme-SPOON (Feinstein et al., 2023; Radica et al., 2023; Radica, 2024). Radica et al. (2023) compared the data to a grid of radiative-convective-thermochemical equilibrium models, while Taylor et al. (2023) performed retrievals with three different frameworks on the spectrum. The results of Radica et al. (2023), including a broadened Na feature and a solar to super-solar metallicity, were largely consistent with those of Nikolov et al. (2018) and Nikolov et al. (2022b). However, they found potential for a hazy terminator from an enhanced aerosol scattering slope at wavelengths below m. This was attributed to a possible degeneracy with the Na abundance, as the NIRISS bandpass contains only a partial Na feature at the blue edge of the spectrum, which could masquerade as a scattering slope (Taylor et al., 2023). Additionally, Taylor et al. (2023) inferred a higher CO2 abundance than predicted by the equilibrium models in Radica et al. (2023). The latter found best-fit equilibrium models with a 1-5x solar metallicity and solar C/O ratio, implying a CO2 abundance of . However, the former retrieved a CO2 abundance more broadly consistent with a 10solar composition (e.g., with Aurora).
@@ 4.1 Revisiting the Atmospheric Composition of WASP-96 b with GP-Aided Retrievals
The retrieved parameter values for each retrieval configuration are summarized in Appendix C, Table 2. Key differences in the retrieved abundances are shown in Figure 6, while Figure 7 shows the posterior distributions for the GP hyperparameters. Overall, the median estimates of the retrieved abundances remain largely consistent across retrieval cases, as expected for well-characterized data. However, for both H2O and CO2—the two most tightly constrained molecules—the precision decreases in the GP-aided cases. In particular, we find that the true CO2 abundance may be difficult to estimate from the partial coverage of its absorption band in the NIRISS bandpass, resulting in a posterior distribution that is 1 dex broader than in the traditional retrieval.
Our new inferred CO2 abundance is consistent with both the super-solar metallicity estimate of Taylor et al. (2023) and the solar metallicity estimate of Radica et al. (2023). While the median estimates for H2O abundance remain consistent between retrieval cases, accounting for underlying correlated noise in the data leads to a wider posterior distribution (an increase of dex). These results imply that the expected precision capability of NIRISS observations on both CO2 and H2O abundance constraints is less than previously predicted with traditional retrievals. The posterior distributions for CO, Na, and K remain largely consistent across the retrieval cases, implying that the traditional retrieval was able to characterize these accurately and with the true expected precision from NIRISS observations.
@@ 4.2 Interpreting GP Hyperparameters in WASP-96 b Retrievals
Across the GP-aided retrievals, there is evidence of a global correlation throughout the entire spectrum with an amplitude of 60-100 ppm and a length scale of 0.01 m. The retrieved amplitude of the global kernel is higher when a local kernel is not considered (85 ppm) than when a local kernel is considered (70-75 ppm). This is likely due to the global kernel compensating for localized unknown features with a higher amplitude which can be characterized by local kernels in the other cases, similar to the behavior seen in retrievals of the synthetic dataset. Performing a model comparison, we calculate a model preference (e.g., ‘detection significance’) for the inclusion of a GP relative to the traditional retrieval, for all scenarios. Specifically, the Global, Global and 1 Local, and Global and 2 Local cases are preferred over a traditional retrieval by 4.8, 3.8, and 3.8, respectively, These results suggest that the inclusion of a GP is moderately to strongly preferred over a traditional retrieval.
@@ 4.3 Assessing the Impact of Background Star Contamination
We include the spectrum with uncorrected background contamination in our analysis (Appendix E), although we find that it does not strongly impact our inferred results. The global kernel is able to marginalize over both the contaminants in the data so that local kernels are not needed at those wavelength ranges. Additionally, the retrieved inferences for atmospheric parameters are consistent between both the corrected and uncorrected data. We therefore conclude that background contamination in the spectrum did not significantly contribute to any biasing effects in the analysis of WASP-96 b.
@@ 5 Conclusions and Discussion
With the advent of JWST, the precision of exoplanet spectroscopy has dramatically increased, making it critical to ensure that modeling assumptions do not introduce systematic biases into atmospheric retrievals. In this work, we incorporated Gaussian Processes (GPs) directly into a transmission spectrum retrieval framework, enabling it to marginalize over structured residuals that arise from model–data mismatches. This approach improves the robustness of inferred atmospheric properties, particularly when dealing with incomplete models or unaccounted-for noise. We summarize our main findings below:
@@
Incorporating Gaussian Processes into retrieval frameworks enables more accurate atmospheric inferences, particularly in the presence of contaminated or imperfect data.
@@
GP-aided retrievals not only improve model fits to the data, but also identify regions of the spectrum where model–data mismatches occur.
@@
We find that the combined global and local GP framework minimizes the bias–variance tradeoff and achieves the lowest mean squared error across most species, providing an interpretable and balanced posterior distribution in the presence of structured residuals.
@@
For H2O, one of the best-constrained molecules in the NIRISS bandpass, our GP-aided retrievals yield a broader posterior (0.1 dex wider) than traditional methods, suggesting that previous estimates may be over-precise.
@@
In the case of CO2, the broader constraints (1 dex) obtained with GP-aided retrievals suggest that discrepancies in the literature may stem from over-precision rather than true disagreement. The difficulty of constraining CO2 from a single, partially covered absorption band in the NIRISS bandpass likely contributes to this effect.
@@
Our GP framework reveals a correlated noise structure across the NIRISS bandpass in WASP-96 b and localizes two regions of significant model–data mismatch, including one overlapping with the CO2 feature at 2.7 m. This mismatch may bias the retrieved CO2 abundance when using traditional retrievals that do not account for correlated noise.
@@ 5.1 Choice of Hyperparameter Models and Priors
Selecting an appropriate GP kernel can be challenging, particularly when the nature of the correlated noise is not well understood a priori. Ideally, the kernel should approximate the physical processes that give rise to the residual structure in the data. For instance, time-series observations of stellar variability have been modeled using damped harmonic oscillator kernels (Foreman-Mackey et al., 2017; Pereira et al., 2019), while similar oscillatory kernels have been employed to mitigate stellar effects in transit light curves (e.g., Radica et al., 2024; Coulombe et al., 2025). In the context of exoplanet light curve fitting, both global RBF and Matérn-3/2 kernels have been used to capture systematic correlations between adjacent observations (e.g., Gibson et al., 2012; Evans et al., 2015; Fortune et al., 2024). Likewise, Czekala et al. (2015) applied both global and local Matérn-3/2 kernels in wavelength space to fit stellar spectral models to observed data.
@@
We adopt a radial basis function (RBF) kernel for both the global and local components. This choice is motivated by its simplicity and its ability to approximate smooth, correlated structures in the data. For the global kernel, the width of the NIRISS point spread function (PSF) implies that adjacent wavelength bins are correlated in a way that decays with inter-pixel distance (Albert et al., 2023), consistent with the type of behavior modeled using Matérn-3/2 kernels in stellar spectra (Czekala et al., 2015).
@@
For local kernels, the RBF is well-suited to approximate unmodeled spectral features in low-resolution data, even without prior knowledge of their exact shape. We compare the performance of the RBF and Matérn-3/2 kernels on our synthetic dataset and find no significant differences in the retrieved atmospheric properties. We therefore adopt the RBF kernel throughout this work. We do not consider periodic kernels in this work as, to the best of our knowledge, no clear wavelength-periodic structure has been identified in NIRISS spectra to date.
@@
We also examine the impact of hyperparameter prior (hyperprior) selection on retrieval outcomes. In Appendix A, we compare log-uniform and uniform priors for the GP hyperparameters using synthetic datasets. As expected, we find that hyperprior choice does not significantly affect the retrieved atmospheric parameters in most cases. This result is consistent with Chen & Wang (2018), who showed that while hyperpriors can influence the best-fit hyperparameters, they have limited impact on the inferred model parameters.
However, we find that hyperpriors do influence the retrieval’s ability to localize features in the spectrum. In particular, uniform priors on both amplitude and length scale lead to better constraints on the central wavelength of local kernels. For this reason, we adopt uniform priors throughout this work. To ensure that our prior ranges do not restrict the inference, we select bounds that encompass any physically plausible contribution to the transmission spectrum. For example, we allow kernel amplitudes to vary from zero (no GP contribution) to 10,000 ppm (equivalent to a 1% transit depth), which is conservatively larger than any expected unknown absorber or contaminant. We recommend that future studies carefully assess the influence of hyperprior choices and adopt physically motivated ranges consistent with the observational setup and scientific goals.
@@ 5.2 Future Directions
Previous work has also explored methods to analytically estimate the spectral covariance matrix directly from the data (e.g., Holmberg & Madhusudhan, 2023; Fortune et al., 2024). While such approaches can characterize noise structure without imposing model assumptions, they do not directly address biases introduced by model deficiencies. However, comparing the analytically derived and retrieval-inferred covariance matrices may help disentangle model–data degeneracies, allowing sources of uncertainty to be more definitively attributed to either data artifacts or modeling limitations. Although this comparison is beyond the scope of the present work, future efforts integrating both methodologies could yield valuable insights into the limitations of current retrieval frameworks.

### v2

@@ 3.2 Recovery of Injected Noise and Feature
In all three cases, we find that the retrieval frameworks are able to provide both precise and accurate estimates of the underlying correlation. Both the global and local kernel amplitudes are consistent with the true amplitudes, with the true values falling within the “single standard deviation” range of the posterior distribution (16th and 84th percentile). The global length scale is similarly constrained in the global-only case, although it is slightly overestimated in the combined case (with the true value falling in the percentile of the posterior distribution, consistent within for a Gaussian distribution). Although the local kernel does overestimate the wavelength and width of the injected feature in both the local-only and combined case, these do not significantly affect its ability to retrieve accurate posterior distributions for atmospheric parameters, and are still within a reasonable margin of error away from the injected values (for both, the true value falls in the percentile of each posterior distribution). Specifically, the retrieved central wavelength is only 0.01 m from the true value, which we attribute to randomness in the noise instance used to simulate the data and the low resolution of the data. The exact central wavelength may be difficult to constrain for a feature characterized by only 5 data points, but 0.01 m is within a single bin width of the true wavelength. As such, we consider this an accurate inference.
Notably, the global-only case retrieves a higher global correlation amplitude than the combined case. This is due to the absence of a local kernel in the global-only case, leading to overcompensation; to account for the 1.65 m feature without a local kernel, the global kernel retrieves a higher amplitude that can marginalize over the feature. Thus, in the combined case, the inferred amplitude is lower, with the true injected value being within of the median inferred amplitude.
@@ 4 Application to the JWST NIRISS/SOSS Spectrum of WASP-96 b
Radica et al. (2023) presented the near-infrared (NIR) transmission spectrum of WASP-96 b observed with JWST NIRISS/SOSS (0.6-2.8 m) as part of the JWST Early Release Observations (ERO) program. The spectrum is split into two overlapping SOSS orders (0.6–1.0 m and 0.85–2.8 m), with an additional third order not considered within the data reduction process. The data was reduced via the exoTEDRF pipeline111Formerly known as supreme-SPOON (Feinstein et al., 2023; Radica et al., 2023; Radica, 2024). Radica et al. (2023) compared the data to a grid of radiative-convective-thermochemical equilibrium models, while Taylor et al. (2023) performed retrievals with three different frameworks on the spectrum. The results of Radica et al. (2023), including a broadened Na feature and a solar to super-solar metallicity, were largely consistent with those of Nikolov et al. (2018) and Nikolov et al. (2022b). However, they found potential for a hazy terminator from an enhanced aerosol scattering slope at wavelengths below m. This was attributed to a possible degeneracy with the Na abundance, as the NIRISS bandpass contains only a partial Na feature at the blue edge of the spectrum, which could masquerade as a scattering slope (Taylor et al., 2023). Additionally, Taylor et al. (2023) inferred a higher CO2 abundance than predicted by the equilibrium models in Radica et al. (2023). The latter found best-fit equilibrium models with a 1–5 solar metallicity and solar C/O ratio, implying a CO2 abundance of . However, the former retrieved a CO2 abundance more broadly consistent with a 10 solar composition (e.g., with Aurora).
@@ 4.1 Revisiting the Atmospheric Composition of WASP-96 b with GP-Aided Retrievals
The retrieved parameter values for each retrieval configuration are summarized in Appendix B, Table 2. Key differences in the retrieved abundances are shown in Figure 6, while Figure 7 shows the posterior distributions for the GP hyperparameters. Overall, the median estimates of the retrieved abundances remain largely consistent across retrieval cases, as expected for well-characterized data. However, for both H2O and CO2 (the two most tightly constrained molecules) the precision decreases in the GP-aided cases. In particular, we find that the true CO2 abundance may be difficult to estimate from the partial coverage of its absorption band in the NIRISS bandpass, resulting in a posterior distribution that is 1 dex broader than in the traditional retrieval. This is consistent with the usage of a GP noise model, which inherently allows for additional flexibility in the model and thus leads to wider posterior distributions for the model parameters.
The inferred CO2 abundance from the GP noise model retrievals is consistent with both the super-solar metallicity estimate of Taylor et al. (2023) and the solar metallicity estimate of Radica et al. (2023). While the median estimates for H2O abundance remain consistent between retrieval cases, accounting for any underlying correlated noise in the data again leads to a wider posterior distribution (an increase of dex). These results imply that the expected precision capability of NIRISS observations on both CO2 and H2O abundance constraints may be less than previously predicted with traditional retrievals. The posterior distributions for CO, Na, and K remain largely consistent across the retrieval cases, implying that the traditional retrieval was able to characterize these accurately and with the true expected precision from NIRISS observations.
@@ 4.2 Interpreting GP Hyperparameters in WASP-96 b Retrievals
Across the GP-aided retrievals, there is evidence of a global correlation throughout the entire spectrum with an amplitude of 60-100 ppm and a length scale of 0.01 m. The retrieved amplitude of the global kernel is higher when a local kernel is not considered (85 ppm) than when a local kernel is considered (70-75 ppm). This is likely due to the global kernel compensating for localized unknown features with a higher amplitude which can be characterized by local kernels in the other cases, similar to the behavior seen in retrievals of the synthetic dataset. Performing a model comparison, we calculate a model preference (e.g., ‘detection significance’) for the inclusion of a GP relative to the traditional retrieval, for all scenarios. Specifically, the Global, Global and 1 Local, and Global and 2 Local cases are preferred over a traditional retrieval by 4.8 (), 3.8 (5.78), and 3.8 (5.79), respectively. These results suggest that the inclusion of a GP is strongly preferred over a traditional retrieval (Trotta, 2008; Benneke & Seager, 2013).
However, a global-only kernel is moderately preferred over either combined kernel (i.e., the summed global and local kernels). Furthermore, the length scale and amplitude posterior distributions for both local kernels are not well-constrained and consistent with zero (Figure 7). These inferences demonstrate that hyperparameter posterior distributions can also diagnose whether a given kernel is needed, while not relying solely on Bayesian model comparisons. In general, the use of more complex kernels with combinations of local and global kernels may therefore be a preferred, less assumptive approach for marginalizing over any possible sub-structure. Unneeded kernels will both be penalized in Bayesian evidence and have hyperparameter posterior distributions consistent with zero.
@@ 4.3 Assessing the Impact of Background Star Contamination
We include the spectrum with uncorrected background contamination in our analysis, although we find that it does not strongly impact our inferred results. The global kernel is able to marginalize over both the contaminants in the data so that local kernels are not needed at those wavelength ranges. Additionally, the retrieved inferences for atmospheric parameters are consistent between both the corrected and uncorrected data. We therefore conclude that background contamination in the spectrum did not significantly contribute to any biasing effects in the analysis of WASP-96 b. The retrieved spectra and posterior distributions for the dataset with uncorrected contamination can be found in the repository for this manuscript222https://osf.io/pksnq/.
@@ 5 Conclusions and Discussion
With the advent of JWST, the precision of exoplanet spectroscopy has dramatically increased, making it critical to ensure that modeling assumptions do not introduce systematic biases into atmospheric retrievals. In this work, we incorporated Gaussian Processes (GPs) directly into a transmission spectrum retrieval framework, enabling it to marginalize over structured residuals that arise from model–data mismatches, including missing absorbers in the model or data unknowns. This approach improves the robustness of inferred atmospheric properties, particularly when dealing with incomplete models or unaccounted-for noise. We summarize our main findings below:
@@
Incorporating Gaussian processes into retrieval frameworks may enable more accurate atmospheric inferences, particularly in the presence of unknowns that can be well-approximated by the chosen GP kernels.
@@
GP-aided retrievals can not only improve model fits to the data, but also identify regions of the spectrum where model–data mismatches occur.
@@
The combined global and local GP framework minimizes the bias–variance tradeoff for synthetic data and achieves the lowest mean squared error across most species, providing an interpretable and balanced posterior distribution in the presence of structured residuals.
@@
For H2O, one of the best-constrained molecules in the NIRISS bandpass, our GP-aided retrievals are preferred by Bayesian evidence and yield a broader posterior (0.1 dex wider) than traditional methods, suggesting that previous estimates of H2O abundance may be over-precise.
@@
In the case of CO2, the broader constraints (1 dex) and higher Bayesian evidence obtained with GP-aided retrievals suggest that current discrepancies in the literature may stem from over-precision rather than true disagreement. The difficulty of constraining CO2 from a single, partially covered absorption band in the NIRISS bandpass likely contributes to this effect.
@@
Our GP framework reveals an underlying correlated noise structure across the NIRISS bandpass in observations of WASP-96 b and localizes two regions of significant model–data mismatch, including one overlapping with the CO2 feature at 2.7 m. This mismatch may bias the retrieved CO2 abundance when using traditional retrievals that do not account for correlated noise.
@@ 5.1 Choice of Hyperparameter Models and Priors
Selecting an appropriate GP kernel can be challenging, particularly when the nature of the correlated noise is not well understood a priori. Ideally, the kernel should approximate the physical processes that give rise to the residual structure in the noise. For instance, time-series observations of stellar variability have been modeled using damped harmonic oscillator kernels (Foreman-Mackey et al., 2017; Pereira et al., 2019), while similar oscillatory kernels have been employed to mitigate stellar effects in transit light curves (e.g., Radica et al., 2024; Coulombe et al., 2025). In the context of exoplanet light curve fitting, both global SE and Matérn-3/2 kernels have been used to capture systematic correlations between adjacent observations (e.g., Gibson et al., 2012; Evans et al., 2015; Fortune et al., 2024). Likewise, Czekala et al. (2015) applied both global and local Matérn-3/2 kernels in wavelength space to fit stellar spectral models to observed data.
@@
We adopt a squared exponential (SE) kernel for both the global and local kernels. This choice is motivated by its simplicity and its ability to approximate smooth, correlated structures in the data. For the global kernel, the width of the NIRISS point spread function (PSF) implies that adjacent wavelength bins are correlated in a way that decays with inter-pixel distance (Albert et al., 2023), consistent with the type of behavior modeled using Matérn-3/2 kernels in stellar spectra (Czekala et al., 2015). In the local case, the chosen structure of the kernel makes it ideal for marginalizing over effects in the spectrum that follow a pseudo-Gaussian shape, such as features from absorbers not accounted for within the model or contaminating stellar features. However, this choice of kernel will likely be unable to account for higher order effects, such as three-dimensional effects for tidally locked planets (e.g., Blecic et al., 2017; Caldas et al., 2019).
@@
For local kernels, the squared exponential (SE) kernel is well-suited to approximate unmodeled spectral features in low-resolution data, even without prior knowledge of their exact shape. We compare the performance of the squared exponential and Matérn-3/2 kernels on our synthetic dataset, but find no significant differences in the retrieved atmospheric properties, and therefore adopt the former throughout this work. We do not consider periodic kernels in this work as, to the best of our knowledge, no clear wavelength-periodic structure has been identified in NIRISS spectra to date. We note that both the Matérn-3/2 and SE kernel are realizations of the generalized Matérn kernel equation. This generalized equation can provide a more flexible kernel realization if the additional hyperparameter , representing the roughness of the covariance, is solved for. However, we consider this outside the scope of this work.
@@
We further examine the impact of hyperparameter prior (hyperprior) selection on retrieval outcomes. We compare log-uniform and uniform priors for the GP hyperparameters using synthetic datasets. As expected, we find that hyperprior choice does not significantly affect the retrieved atmospheric parameters in most cases. This result is consistent with Chen & Wang (2018), who showed that while hyperpriors can influence the best-fit hyperparameters, they have limited impact on the inferred model parameters.
However, we find that hyperprior selection does influence how well the retrieval can constrain information about the correlation in the data. In particular, uniform priors on kernel amplitude and length scale may yield narrower posteriors on the location and strength of local correlations, particularly in low-SNR regimes. In contrast, log-uniform priors allocate greater density to small values and may lead to broader posteriors centered near negligible kernel contributions. This behavior reflects the interplay between the prior volume and the informativeness of the likelihood: when the data only weakly constrain the kernel, the choice of prior becomes non-negligible. We adopt uniform priors throughout to more readily detect and quantify non-zero kernel structures, acknowledging that this choice may bias against the null (no-correlation) hypothesis. To ensure that our prior ranges do not restrict the inference, we select bounds that encompass any physically plausible contribution to the transmission spectrum. For example, we allow kernel amplitudes to vary from zero (no GP contribution) to 10,000 ppm (equivalent to a 1% transit depth), which is conservatively larger than any expected unknown absorber or contaminant.
We also find that the kernel amplitude and length scale become partially degenerate in two limiting regimes. When the amplitude approaches zero, the kernel contribution vanishes and the likelihood becomes insensitive to the length scale, rendering it unconstrained. Conversely, when the length scale becomes smaller than the typical bin spacing of the data (m), the kernel contribution becomes sharply localized, effectively acting as a spike at a single data point. In this limit, the kernel can mimic additional white noise, which could introduce redundancy with the noise model and limit its identifiability.
The choice of hyperparameter priors can also strongly impact the relative model preference (Bayes factor) of a GP noise model when compared to a traditional white noise model against the data. Bayes factors are the ratio of the Bayesian evidences of two models, each of which can be thought of as the weighted average of the Bayesian likelihood across the prior volume. Their ratio corresponds to the statistical significance of the preference of one model over another (e.g., Trotta, 2008). By expanding the prior range of the hyperparameters, one can thus inadvertently dilute the Bayesian evidence of the GP noise model even in the case of correlated noise in the data. This effect can lead to selectively preferring a traditional noise model, which will in turn require a more complex physical model to overcome the correlated noise structure without a GP. A balance must therefore be struck between a broad hyperprior that contains the possible hyperparameter values to avoid biasing the kernel inference, and one that is not overly broad so as to dilute the evidence (for further discussion of this effect, see, e.g., MacKay, 2003; Trotta, 2008; Llorente et al., 2023). We consider further exploration of prior selection beyond the scope of this work, but recommend that future studies carefully assess the influence of hyperprior choices and adopt physically motivated ranges consistent with the observational setup and scientific goals.
@@ 5.2 Future Directions
Previous work has also explored methods to analytically estimate the spectral covariance matrix directly from the data (e.g., Holmberg & Madhusudhan, 2023; Fortune et al., 2024). While such approaches can characterize noise structure without imposing model assumptions, they are mathematically complex, and do not directly address biases introduced by model deficiencies. However, comparing the analytically derived and retrieval-inferred covariance matrices may help disentangle model–data degeneracies. Additionally, the relative simplicity of the inferred covariance matrix from the retrieval allows for scalability to multiple spectra, providing an opportunity for identification of trends from instrumental or physical effects. Although a comparison between the methods is beyond the scope of the present work, future efforts integrating both methodologies could yield valuable insights into the limitations of current retrieval frameworks.